In loading a delivery van a driver pushed a 15 kg crate for 5 metres up a ramp. He had to push with a force of 132N and the vertical height gained was 1.3m. What was the efficiency of pushing this crate up the ramp onto the back of the van
1
Expert's answer
2020-01-29T13:54:07-0500

length of the ramp ="\sqrt{(5^2+(1.3)^2)}=\sqrt{(25+1.69)}=5.17m"

efficiency= "\dfrac{net \ work \ done }{total\ work \ done}"


net work done is equal to the change in the potential energy of the box is

"mgh" "=15\times10\times1.3=195 / joules"

total work done is equal to work done by the driver up the plank is

"F\times d" "=132\times5.17=682.44 joules"

"efficiency=\dfrac{ 195}{682.44}=0.29"

"efficiency % =" %= "0.29\times100= 29" %